OpenStudy (anonymous):

ln(1+e^-x) =3 solve for x ???? cant figure these out

OpenStudy (jamesj):

Take exp first of both sides, then 1 + e^-x = e^3 Now can you solve that for x?

OpenStudy (anonymous):

start by writing in equivalent logarithmic form to get \[1+e^{-x}=e^3\] and then oh again. man is the site slow

OpenStudy (anonymous):

is it e^-x = e^3 - 1 then -x = 3-1 then -x = 2 then x = -2 sorry if thats hard to follow

OpenStudy (jamesj):

No definitely not, because how did the -1 get into the index?

OpenStudy (jamesj):

\[1 + e^{-x} = e^3 \implies e^{-x} = e^3 - 1\]

OpenStudy (jamesj):

Now given that, how do you find x?

OpenStudy (jamesj):

If I told you in general that \[e^x = a\] what is x?

OpenStudy (anonymous):

is it one i am honestly not sure

OpenStudy (jamesj):

ln x is the inverse function of e^x Hence if e^x = a, then x = ln(a)

OpenStudy (jamesj):

That is why when we started the problem we said that \[\ln(1+e^-x) =3 \implies 1 + e^{-x} = e^3\]

OpenStudy (jamesj):

make that: \[\ln(1+e^{-x}) =3 \implies 1 + e^{-x} = e^3\]

OpenStudy (jamesj):

because in general \[\ln(a) = b \iff a = e^b\]

OpenStudy (jamesj):

So now we have \[e^{-x} = e^3 - 1\] What do you get if you take the ln of both sides?

OpenStudy (jamesj):

What is \[\ln(e^{-x})\]?

OpenStudy (jamesj):

and you're not even here. Alright. Good luck.

OpenStudy (anonymous):

i thought that the e's were just eliminated but i am waiting on my gf to bring my textbook. looks like i have some serious studying to do

OpenStudy (jamesj):

For the record, what is \[\ln(e^2)\] ?